OpenStudy (anonymous):

(x+6)/12 divided by (x-8)/10 A.5x/x-8 B. 5x/2(x-4) C.5(x+6)/6(x-8) D.(x+6)(x-8)/30

OpenStudy (anonymous):

c

OpenStudy (anonymous):

= x + 6 * 10 ----- -- 12 x-8 = 5(x+6) / 6(x-8)

OpenStudy (anonymous):

yea C

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

no problems kaira....

OpenStudy (anonymous):

yw